# High Mass Dark Matter Searches With the High Speed Flux From the Large Magellanic Cloud

### v1

@@ 1 Introduction
For decades, astrophysicists have posited the existence of dark matter to explain several astronomical observations, including galactic rotation curves, the cosmic microwave background, gravitational lensing observations and galactic structure formation. It has long been appreciated that the unit mass of dark matter in the galactic halo can be orders of magnitude larger than Standard Model particles [Witten:1984rs, Farhi:1984qu, DeRujula:1984axn, Goodman:1984dc, Drukier:1986tm, Nussinov:1985xr, Bagnasco:1993st, Alves:2009nf, Kribs:2009fy]. Such massive states can be formed through dark sector nucleosynthesis [Krnjaic:2014xza, Wise:2014jva, Gresham:2017cvl, Acevedo:2020avd], dissipative processes in the early universe [Buckley:2017ttd, Chang:2018bgx, Bramante:2023ddr, Bramante:2024pyc, Lu:2024xnb], or a first order phase transition [Witten:1984rs, Bai:2018dxf]. Due to their low number density, such heavy halo constituents can have relatively high cross sections with Standard Model particles without conflicting with known bounds, and there are simple models that imply large cross sections for asymmetric composites via a weak coupling, for example a Higgs portal coupling to nucleons [Acevedo:2021kly, Acevedo:2024lyr, Bleau:2025klr]. At large cross sections, dark matter can scatter multiple times with normal matter as it passes through the detector and overburden material. This means incoming dark matter may lose substantial energy to the overburden before reaching an underground detector, motivating searches nearer to the surface and even in outer space. Fortuitously, there are no known Standard Model background processes that leave a signature similar to that of such heavy dark matter, which deposits many nuclear recoils along a very straight trajectory. Because many solid-state multiscatter detectors require a minimum recoil energy per unit path length to register successive events, their sensitivity is especially dependent on the dark matter velocity distribution. In all prior work, multiscatter searches have assumed dark matter follows the Standard Halo Model (SHM).
@@
The SHM [PhysRevD.33.3495] is the simplest and most commonly adopted model for the dark matter distribution in the Galaxy. In the SHM, the dark matter halo is assumed to be isothermal and spherical, with an isotropic Maxwell-Boltzmann velocity distribution which has a peak speed equal to the local circular speed. However, the actual dark matter distribution may have significant differences with respect to the SHM. In particular, high-resolution cosmological simulations that include both dark matter and baryons indicate that, while a Maxwellian distribution provides a reasonable fit to the local dark matter velocity distribution in simulated Milky Way (MW) analogues, the results carry large uncertainties [Bozorgnia:2016ogo, Kelso:2016qqj, Sloane:2016kyi, Bozorgnia:2017brl, Bozorgnia:2019mjk, Poole-McKenzie:2020dbo, Lawrence:2022niq, Kuhlen:2013tra, Lacroix:2020lhn, Santos-Santos:2023ubx, Herrera:2023fpq, Herrera:2021puj].
@@
One of the most important deviations from the SHM originates from the impact of the Large Magellanic Cloud (LMC) on the local dark matter distribution. The LMC is the most massive satellite and the most recent merger of the MW. It has recently completed its first pericentric passage and is currently at a distance of $\sim 50$ kpc, moving at a speed of $321\pm 24$ km/s with respect to the MW [Besla:2007kf, Kallivayalil:2013xb, Patel:2017, Pietrzynski:2019]. Recent studies employing both cosmological [Smith-Orlik:2023kyl, Reynoso-Cordova:2024xqz] and idealized [Besla:2019xbx, Donaldson:2021byu] simulations demonstrate that the LMC significantly boosts the local dark matter velocity distribution to velocities above the Milky Way’s escape speed. This boost stems from the presence of high-speed dark matter particles originating from the LMC, as well as the gravitational influence of the LMC accelerating MW dark matter particles [Besla:2019xbx, Smith-Orlik:2023kyl]. As a result, the LMC can shift direct detection exclusion limits by several orders of magnitude toward smaller dark matter-nucleon cross sections and lower dark matter masses, with the most pronounced impact for low-mass dark matter particles [Smith-Orlik:2023kyl, Reynoso-Cordova:2024xqz].
@@
Previous work [Jacobs:2014yca, Bhoonah2021] computed exclusion limits for heavy dark matter, where [Bhoonah2021] particularly used the SHM in its analysis and identified two existing searches for massive particles, namely the Skylab cosmic ray search and a search for monopoles at the Ohya quarry. The remainder of this paper is structured as follows. Section 2 describes the simulated MW-LMC analogue used in this work and our procedure for selecting a Solar neighborhood consistent with the observed LMC-Sun geometry. In Section 3 we lay out the multiscatter detector response formalism, including interaction models (spin-independent nucleon and contact cross sections), energy loss in overburden, directional efficiencies, and in Section 4 we use discrete simulation samples to compute fluxes and energy thresholds for Ohya and Skylab. We then present the resultant limits in Section 5, comparing MW-LMC predictions to the SHM, and compare these to other searches. We conclude by summarizing the impact of the LMC-induced high-speed velocity tail on heavy-dark matter searches, discussing latitude-dependent flux effects, and outlining applications to future ground and orbital multiscatter detectors.
@@ 2.1 MW-LMC analogue
We use a MW-LMC analogue from the Auriga magneto-hydrodynamical simulations [Grand:2016mgo, Grand:2024]. The Auriga project [Grand:2016mgo] contains a set of cosmological zoom-in simulations of isolated MW-mass halos, selected from a (100 Mpc)3 periodic cube (L100N1504) from the EAGLE project [Schaye:2014tpa, Crain:2015poa]. The simulations were carried out using the moving-mesh code Arepo [Springel:2009aa] and implement a galaxy formation subgrid model which includes star formation, supernova feedback, active galactic nuclei, black hole formation, metal cooling, and background UV/X-ray photoionisation radiation [Grand:2016mgo]. The simulations adopt the Planck-2015 [Planck:2015fie] cosmological parameters: $\Omega_{m}=0.307$, $\Omega_{\rm bar}=0.048$, $H_{0}=67.77\penalty 10000\ {\rm km\penalty 10000\ s^{-1}\penalty 10000\ Mpc^{-1}}$. We use the standard resolution level (Level 4) of the simulations, in which the dark matter particle mass is $m_{\rm DM}\sim 3\times 10^{5}\penalty 10000\ {\rm M}_{\odot}$, the baryonic particle mass is $m_{b}=5\times 10^{4}\penalty 10000\ {\rm M}_{\odot}$, and the Plummer equivalent gravitational softening length is $\epsilon=370$ pc [Power:2002sw, Jenkins2013]. The Auriga simulations reproduce the observed stellar masses, sizes, rotation curves, star formation rates and metallicities of present day MW-mass galaxies.
The LMC’s first pericenter approach occurred $\sim$ 50 Myr ago [Besla:2007kf] at a distance of $\sim 48$ kpc [Besla:2007kf], and the present-day stellar mass of the LMC from observations is $\sim 2.7\times 10^{9}$ M [vanderMarel:2002kq]. We use a simulated MW-LMC analogue whose properties match those of the observed system. The analogue is the re-simulated halo 13 in ref. [Smith-Orlik:2023kyl], corresponding to the Auriga 25 MW halo and its LMC analogue. This system was selected from a sample of 15 MW-LMC analogues in ref. [Smith-Orlik:2023kyl], which were identified based on the present-day stellar mass of the LMC and its distance from host at first pericenter. Due to the large average time ($\sim 150$ Myr) between simulation snapshots in the 15 MW-LMC analogues, it is difficult to accurately identify their present-day snapshot. Therefore, one system (halo 13) was chosen and re-simulated in ref. [Smith-Orlik:2023kyl], such that the average time between snapshots close to the LMC’s pericenter approach is $\sim 10$ Myr, and the present-day snapshot closely matches observations. For this system, the virial mass of the MW host halo is $1.2\times 10^{12}\penalty 10000\ {\rm M}_{\odot}$ and the halo mass of the LMC analogue at infall is $3.2\times 10^{11}\penalty 10000\ {\rm M}_{\odot}$. More details on this system are provided in ref. [Smith-Orlik:2023kyl].
@@ 2.2 Solar region
There is no Solar position in the simulations a priori, and this needs to be specified. As discussed in ref. [Smith-Orlik:2023kyl], we choose the Sun’s position and velocity in the simulated halo such that it matches the observed Sun-LMC geometry. This matching is performed in three steps. In the first step, we determine the orientations of the stellar disk in the MW analogue that reproduce the observed angle with the orbital plane of the LMC analogue. In the second step, for each allowed disk orientation we determine the Sun’s position by requiring that the angles between the LMC’s orbital angular momentum and the Sun’s position and velocity vectors in the simulations match their observed values. In the third step, we determine the best fit Sun’s position as the one that yields the closest match of the angles between the Sun’s velocity vector and the LMC’s position and velocities with their observed values.
The Solar region is defined as the overlap of two regions: a spherical shell with radii between 6 to 10 kpc from the MW analogue’s center (with the Sun at a galactocentric distance of $\sim 8$ kpc), and a cone with its axis aligned with the best fit Sun’s position, vertex at the galactic center, and an opening angle of $\pi/4$ radians. As detailed in ref. [Smith-Orlik:2023kyl], the size of the Solar region is chosen such that it would be sensitive to the best fit Sun’s position and also include several thousand dark matter particles.
@@ 3 Heavy Dark Matter and Plastic Etch Detectors
We now lay out our treatment of heavy dark matter scattering in solid state detectors, and how this treatment is adapted to account for the high velocity component of dark matter assocaited with the LMC. Following [Bhoonah2021], we consider two different models of dark matter interactions with nuclei. The first is a spin-independent nuclear scattering cross section. In this case, omitting nuclear form factors, the dark matter-nuclear cross section is defined as
@@
where $\rho_{\chi}=0.3\ \mathrm{GeV}/\mathrm{cm}^{3}$ is the local dark matter density [Eilers_2019dmspeed], $m_{\chi}$ is the dark matter mass, $A_{d}besttime$ is the area of the detector, and $T$ is the observation period for the detector. $f(\mathbf{v})$ is the local dark matter velocity distribution in the reference frame of the earth and $I(\mathbf{v})$ is a time average over detector quantities,
@@
We now describe the bounds, which are computed using the same method as [Bhoonah2021]. The experiments detected no straight line tracks through the detectors consistent with dark matter. To be conservative, we assume that the Ohya and Skylab experiments were unlucky and that tracks due to dark matter would have been observed 90% of the time. This sets the average number of observed dark matter particles to be $N_{c}=-\ln(0.1)=2.3$, assuming Poisson statistics. The maximum dark matter mass an experiment excludes can be found by matching the number of observed dark matter particles, $N_{c}$, with the total expected number passing through the detector, $N(m_{\chi})$. Next, to compute upper and lower bounds on the cross section, we determine the dark matter cutoff speed $v_{0}$ above which one sees $N_{c}$ particles pass through the detector, though not necessarily triggering it. We do so by solving,
@@
Here, $\tilde{n}_{D}$ is the effective number density of the detector, given in Equation 9 or 10, and $x_{D}$ is length of the dark matter particle’s path through the detector. $\tilde{n}_{S}$ and $x_{S}$ are defined similarly for the solid overburden. $\tilde{\mathcal{N}}_{A}=\int\tilde{n}_{A}dx$ is the effective column density of the atmosphere along the particles path. To be conservative, we assume the dark matter will traverse the overburden and hit the detector at the largest possible angle: the cutoff angle $\theta_{\mathrm{tol}}$. Also in Equation 13, $\sigma$ denotes either the per-nucleon cross section $\sigma_{\chi n}$ or the contact cross section $\sigma_{C}$.
@@ 4 Modelling the Flux
$\hat{v}=-\begin{bmatrix}v\cos(\delta_{\chi})\cos(\alpha_{\chi})\\ v\cos(\delta_{\chi})\sin(\alpha_{\chi})\\ v\sin(\delta_{\chi}),\end{bmatrix}$ (18)
where $\delta_{\chi}$ and $\alpha_{\chi}$ are the declination and right ascension of the incoming dark matter in the sky.
In this case, the integral in Equation 16 can be calculated analytically,
@@ 5 Improved Bounds from Skylab and Ohya
In this section, we describe the improvements to the dark matter exclusion plots from plastic etch detectors at the Ohya quarry and onboard the Skylab satellite. The plastic etch detector at Ohya searched for magnetic monopoles. The other was a cosmic ray search for highly charged nuclei aboard the ISS’s predecessor satellite Skylab. Both experiments used detectors made of sheets of plastic polymer. When a particle passes through the plastic, it can damage the molecular bonds. This damage becomes visible as a hole after an acid wash. [Bhoonah2021] treated a hole burrowed entirely through the detector as the dark matter detection signal, which neither search observed. We present the relevant parameters from the experiments in Table 1.
@@
Column Number Density of
Atmospheric Overburden
$960\>\frac{\text{g}}{\text{cm}^{2}}$ $0\>\frac{\text{g}}{\text{cm}^{2}}$
@@
Table 1: The experimental parameters from the Skylab and Ohya experiments relevant to this work. These are taken from [Bhoonah2021].
The Ohya detector ran for 2.1 years. It was made of Lexan $\mathrm{(C_{16}H_{14}O_{3}})$ with an energy threshold of $0.3\>\text{GeV}/\text{cm}$, a density of $1.2\>\text{g}/\text{cm}^{3}$ and an effective area of $2442\>\text{m}^{2}$. The detector had a tolerance angle of $18.4\degree$ and a thickness of $0.66\>\text{cm}$ at that tolerance angle. Ohya’s overburden consists of the entire atmosphere, then crust with a thickness of $39$ m at the tolerance angle. Information on the crustal abundances of elements was taken from the Mendeleev package in Python [mendeleev2014], which obtained its data from [haynes2014crc]. For Ohya’s atmospheric overburden, we use the isothermal model, with atmospheric mass density, $\rho_{A}$, exponentially dependent on altitude $h$ [Cappiello:2023],
@@
where $h_{0}=7\>\text{km}$ is a characteristic height and $\rho_{0}=1.3\times 10^{-3}\>\text{g}/\text{cm}^{3}$ is the atmospheric density at ground level. Since the tolerance angle of the detector, $\theta_{\mathrm{tol}}$, is not close to $90\degree$, the column number density of dark matter entering at the tolerance angle is $h\rho_{0}\sec(\theta_{\mathrm{tol}})$, to a very good approximation. We also assume that the elemental ratios of the atmosphere are independent of height with values from [Bramante2023Atm], fixed at 75.6% nitrogen, 23.18% oxygen, and 1.2% argon by mass.
@@
To compute the effective number densities in Equations 9 and 10, we sum over all isotopes of all elements in the material. For the isotope data, we again use the Mendeleev package, which uses isotopic abundance data from [Kondev_2021] and isotopic mass data from [isotopeMasses].
@@
We proceed with a brief discussion on the origin of the exclusion lines and the effect of the LMC on the bounds. At the high mass end, the bound stops because the flux is too small to trigger the detector. The faster particles due to the impact of the LMC increase the flux, improving the bound. For higher cross sections, the particle loses too much energy in the overburden and can’t trigger the detector. Since the loss in the overburden is exponential (see Equation 12), the faster particles due to the LMC have little effect on the bound. At small cross sections, the bound breaks down because the interaction is too weak to trigger the detector. Here, the effect of the increased flux due to the LMC is the most significant. The effect is most pronounced at low masses, where high speed particles more frequently pass through the detector.
@@
Figure 6 shows the result of including the effect of the LMC on Skylab and Ohya bounds when compared to leading bounds for the per-nucleon and contact cases.

### v2

@@ 1 Introduction
For decades, astrophysicists have posited the existence of dark matter to explain several astronomical observations, including galactic rotation curves, the cosmic microwave background, gravitational lensing observations and galactic structure formation. It has long been appreciated that the unit mass of dark matter in the galactic halo can be orders of magnitude larger than Standard Model particles [63, 28, 21, 29, 25, 48, 6, 5, 41]. Such massive states can be formed through dark sector nucleosynthesis [42, 62, 32, 2], dissipative processes in the early universe [18, 19, 16, 15, 46], or a first order phase transition [63, 7]. Due to their low number density, such heavy halo constituents can have relatively high cross sections with Standard Model particles without conflicting with known bounds, and there are simple models that imply large cross sections for asymmetric composites via a weak coupling, for example a Higgs portal coupling to nucleons [3, 1, 11]. At large cross sections, dark matter can scatter multiple times with normal matter as it passes through the detector and overburden material. This means incoming dark matter may lose substantial energy to the overburden before reaching an underground detector, motivating searches nearer to the surface and even in outer space. Fortuitously, there are no known Standard Model background processes that leave a signature similar to that of such heavy dark matter, which deposits many nuclear recoils along a very straight trajectory. Because many solid-state multiscatter detectors require a minimum recoil energy per unit path length to register successive events, their sensitivity is especially dependent on the dark matter velocity distribution. In all prior work, multiscatter searches have assumed dark matter follows the Standard Halo Model (SHM).
@@
The SHM [26] is the simplest and most commonly adopted model for the dark matter distribution in the Galaxy. In the SHM, the dark matter halo is assumed to be isothermal and spherical, with an isotropic Maxwell-Boltzmann velocity distribution which has a peak speed equal to the local circular speed. However, the actual dark matter distribution may have significant differences with respect to the SHM. In particular, high-resolution cosmological simulations that include both dark matter and baryons indicate that, while a Maxwellian distribution provides a reasonable fit to the local dark matter velocity distribution in simulated Milky Way (MW) analogues, the results carry large uncertainties [13, 39, 58, 12, 14, 52, 45, 43, 44, 56, 34, 35].
@@
One of the most important deviations from the SHM originates from the impact of the Large Magellanic Cloud (LMC) on the local dark matter distribution. The LMC is the most massive satellite and the most recent merger of the MW. It has recently completed its first pericentric passage and is currently at a distance of $\sim 50$ kpc, moving at a speed of $321\pm 24$ km/s with respect to the MW [8, 38, 49, 51]. Recent studies employing both cosmological [59, 55] and idealized [9, 24] simulations demonstrate that the LMC significantly boosts the local dark matter velocity distribution to velocities above the Milky Way’s escape speed. This boost stems from the presence of high-speed dark matter particles originating from the LMC, as well as the gravitational influence of the LMC accelerating MW dark matter particles [9, 59]. As a result, the LMC can shift direct detection exclusion limits by several orders of magnitude toward smaller dark matter-nucleon cross sections and lower dark matter masses, with the most pronounced impact for low-mass dark matter particles [59, 55].
@@
Previous work [36, 10] computed exclusion limits for heavy dark matter, where [10] particularly used the SHM in its analysis and identified two existing searches for massive particles, namely the Skylab cosmic ray search and a search for monopoles at the Ohya quarry. The remainder of this paper is structured as follows. Section 2 describes the simulated MW-LMC analogue used in this work and our procedure for selecting a Solar neighborhood consistent with the observed LMC-Sun geometry. In Section 3 we lay out the multiscatter detector response formalism, including interaction models (spin-independent nucleon and contact cross sections), energy loss in overburden, directional efficiencies, and in Section 4 we use discrete simulation samples to compute fluxes and energy thresholds for Ohya and Skylab. We then present the resultant limits in Section 5, comparing MW-LMC predictions to the SHM, and compare these to other searches. We conclude by summarizing the impact of the LMC-induced high-speed velocity tail on heavy-dark matter searches, discussing latitude-dependent flux effects, and outlining applications to future ground and orbital multiscatter detectors.
@@ 2.1 MW-LMC analogue
We use a MW-LMC analogue from the Auriga magneto-hydrodynamical simulations [31, 30]. The Auriga project [31] contains a set of cosmological zoom-in simulations of isolated MW-mass halos, selected from a (100 Mpc)3 periodic cube (L100N1504) from the EAGLE project [57, 20]. The simulations were carried out using the moving-mesh code Arepo [60] and implement a galaxy formation subgrid model which includes star formation, supernova feedback, active galactic nuclei, black hole formation, metal cooling, and background UV/X-ray photoionisation radiation [31]. The simulations adopt the Planck-2015 [4] cosmological parameters: $\Omega_{m}=0.307$, $\Omega_{\rm bar}=0.048$, $H_{0}=67.77\penalty 10000\ {\rm km\penalty 10000\ s^{-1}\penalty 10000\ Mpc^{-1}}$. We use the standard resolution level (Level 4) of the simulations, in which the dark matter particle mass is $m_{\rm DM}\sim 3\times 10^{5}\penalty 10000\ {\rm M}_{\odot}$, the baryonic particle mass is $m_{b}=5\times 10^{4}\penalty 10000\ {\rm M}_{\odot}$, and the Plummer equivalent gravitational softening length is $\epsilon=370$ pc [53, 37]. The Auriga simulations reproduce the observed stellar masses, sizes, rotation curves, star formation rates and metallicities of present day MW-mass galaxies.
The LMC’s first pericenter approach occurred $\sim$ 50 Myr ago [8] at a distance of $\sim 48$ kpc [8], and the present-day stellar mass of the LMC from observations is $\sim 2.7\times 10^{9}$ M [61]. We use a simulated MW-LMC analogue whose properties match those of the observed system. The analogue is the re-simulated halo 13 in ref. [59], corresponding to the Auriga 25 MW halo and its LMC analogue. This system was selected from a sample of 15 MW-LMC analogues in ref. [59], which were identified based on the present-day stellar mass of the LMC and its distance from host at first pericenter. Due to the large average time ($\sim 150$ Myr) between simulation snapshots in the 15 MW-LMC analogues, it is difficult to accurately identify their present-day snapshot. Therefore, one system (halo 13) was chosen and re-simulated in ref. [59], such that the average time between snapshots close to the LMC’s pericenter approach is $\sim 10$ Myr, and the present-day snapshot closely matches observations. For this system, the virial mass of the MW host halo is $1.2\times 10^{12}\penalty 10000\ {\rm M}_{\odot}$ and the halo mass of the LMC analogue at infall is $3.2\times 10^{11}\penalty 10000\ {\rm M}_{\odot}$. More details on this system are provided in ref. [59]. Finally, we note that we use galactocentric coordinates defined such that the DM wind will come from +90 degrees in longitude [13, 59]. We find our directional component agrees with prior work, which sometimes uses the opposite convention for the galactocentric longitudinal coordinates [24].
@@ 2.2 Solar region
There is no Solar position in the simulations a priori, and this needs to be specified. As discussed in ref. [59], we choose the Sun’s position and velocity in the simulated halo such that it matches the observed Sun-LMC geometry. This matching is performed in three steps. In the first step, we determine the orientations of the stellar disk in the MW analogue that reproduce the observed angle with the orbital plane of the LMC analogue. In the second step, for each allowed disk orientation we determine the Sun’s position by requiring that the angles between the LMC’s orbital angular momentum and the Sun’s position and velocity vectors in the simulations match their observed values. In the third step, we determine the best fit Sun’s position as the one that yields the closest match of the angles between the Sun’s velocity vector and the LMC’s position and velocities with their observed values.
The Solar region is defined as the overlap of two regions: a spherical shell with radii between 6 to 10 kpc from the MW analogue’s center (with the Sun at a galactocentric distance of $\sim 8$ kpc), and a cone with its axis aligned with the best fit Sun’s position, vertex at the galactic center, and an opening angle of $\pi/4$ radians. As detailed in ref. [59], the size of the Solar region is chosen such that it would be sensitive to the best fit Sun’s position and also include several thousand dark matter particles.
@@ 3 Heavy Dark Matter and Plastic Etch Detectors
We now lay out our treatment of heavy dark matter scattering in solid state detectors, and how this treatment is adapted to account for the high velocity component of dark matter associated with the LMC. Following [10], we consider two different models of dark matter interactions with nuclei. The first is a spin-independent nuclear scattering cross section. In this case, omitting nuclear form factors, the dark matter-nuclear cross section is defined as
@@
where $\rho_{\chi}=0.3\ \mathrm{GeV}/\mathrm{cm}^{3}$ is the local dark matter density [27], $m_{\chi}$ is the dark matter mass, $A_{d}$ is the area of the detector, and $T$ is the observation period for the detector. $f(\mathbf{v})$ is the local dark matter velocity distribution in the reference frame of the earth and $I(\mathbf{v})$ is a time average over detector quantities,
@@
We now describe the bounds, which are computed using the same method as [10]. The experiments detected no straight line tracks through the detectors consistent with dark matter. To be conservative, we assume that the Ohya and Skylab experiments were unlucky and that tracks due to dark matter would have been observed 90% of the time. This sets the average number of observed dark matter particles to be $N_{c}=-\ln(0.1)=2.3$, assuming Poisson statistics. The maximum dark matter mass an experiment excludes can be found by matching the number of observed dark matter particles, $N_{c}$, with the total expected number passing through the detector, $N(m_{\chi})$. Next, to compute upper and lower bounds on the cross section, we determine the dark matter cutoff speed $v_{0}$ above which one sees $N_{c}$ particles pass through the detector, though not necessarily triggering it. We do so by solving,
@@
Here, $\tilde{n}_{D}$ is the effective number density of the detector, given in Equation 9 or 10, and $x_{D}$ is length of the dark matter particle’s path through the detector. $\tilde{n}_{S}$ and $x_{S}$ are defined similarly for the solid overburden. $\tilde{\mathcal{N}}_{A}=\int\tilde{n}_{A}dx$ is the effective column density of the atmosphere along the particles path. To be conservative, we assume the dark matter will traverse the overburden and hit the detector at the largest possible angle: the cutoff angle $\theta_{\mathrm{tol}}$. In Equation 13, $\sigma$ denotes either the per-nucleon cross section $\sigma_{\chi n}$ or the contact cross section $\sigma_{C}$.
@@ 4 Modelling the Flux
$\hat{v}=-\begin{bmatrix}v\cos(\delta_{\chi})\cos(\alpha_{\chi})\\ v\cos(\delta_{\chi})\sin(\alpha_{\chi})\\ v\sin(\delta_{\chi})\end{bmatrix}$, (18)
where $\delta_{\chi}$ and $\alpha_{\chi}$ are the declination and right ascension of the incoming dark matter in the sky. In this case, the integral in Equation 16 can be calculated analytically,
@@ 5 Improved Bounds from Skylab and Ohya
In this section, we describe the improvements to the dark matter exclusion plots from plastic etch detectors at the Ohya quarry and onboard the Skylab satellite. The plastic etch detector at Ohya searched for magnetic monopoles. The other was a cosmic ray search for highly charged nuclei aboard the ISS’s predecessor Skylab. Both experiments used detectors made of sheets of plastic polymer. When a particle passes through the plastic, it can damage the molecular bonds. This damage becomes visible as a hole after an acid wash. [10] treated a hole burrowed entirely through the detector in a straight line as the dark matter detection signal, which neither search observed. We present the relevant parameters from the experiments in Table 1.
@@
Column Density of Atmospheric Overburden at $\theta_{\mathrm{tol}}$ $960\>\frac{\text{g}}{\text{cm}^{2}}$ $0\>\frac{\text{g}}{\text{cm}^{2}}$
@@
Table 1: The experimental parameters from the Skylab and Ohya experiments relevant to this work. These are taken from [10].
The Ohya detector ran for 2.1 years. It was made of Lexan $\mathrm{(C_{16}H_{14}O_{3}})$ with an energy threshold of $0.3\>\text{GeV}/\text{cm}$, a density of $1.2\>\text{g}/\text{cm}^{3}$ and an effective area of $2442\>\text{m}^{2}$. The detector had a tolerance angle of $18.4\degree$ and a thickness of $0.66\>\text{cm}$ at that tolerance angle. Ohya’s overburden consists of the entire atmosphere, then crust with a thickness of $39$ m at the tolerance angle. Information on the crustal abundances of elements was taken from the Mendeleev package in Python [47], which obtained its data from [33]. For Ohya’s atmospheric overburden, we use the isothermal model, with atmospheric mass density, $\rho_{A}$, exponentially dependent on altitude $h$ [23],
@@
where $h_{0}=7\>\text{km}$ is a characteristic height and $\rho_{0}=1.3\times 10^{-3}\>\text{g}/\text{cm}^{3}$ is the atmospheric density at ground level. Since the tolerance angle of the detector, $\theta_{\mathrm{tol}}$, is not close to $90\degree$, the column number density of dark matter entering at the tolerance angle is $h\rho_{0}\sec(\theta_{\mathrm{tol}})$ to a very good approximation. We also assume that the elemental ratios of the atmosphere are independent of height with values from [17], fixed at 75.6% nitrogen, 23.18% oxygen, and 1.2% argon by mass.
@@
To compute the effective number densities in Equations 9 and 10, we sum over all isotopes of all elements in the material. For the isotope data, we again use the Mendeleev package, which uses isotopic abundance data from [40] and isotopic mass data from [54].
@@
We proceed with a brief discussion on the origin of the exclusion lines and the effect of the LMC on the bounds. At the high mass end, the bound stops because the flux is too small to trigger the detector. The faster particles due to the impact of the LMC increase the flux, improving the bound. For higher cross sections, the particle loses too much energy in the overburden and can’t trigger the detector. Since the loss in the overburden is exponential (see Equation 12), the faster particles due to the LMC have little effect on the bound. At small cross sections, the bound breaks down because the interaction is too weak to trigger the detector. Here, the effect of the increased flux due to the LMC is the most significant. The effect is most pronounced at low masses, where high speed particles more frequently pass through the detector. Figure 6 shows the result of including the effect of the LMC on Skylab and Ohya bounds when compared to leading bounds for the per-nucleon and contact cases.